# Jianxiong Gao, Yichang Liu, Baofeng Yang, Jianfeng Feng, Yanwei Fu

### v1

@@ 4.1 Multi-Modal Fusion Encoder
Inspired by the success of contrastive learning [36] in neuro-decoding applications [7, 27, 14] and multi-modal tasks, and considering that fMRI and EEG signals significantly differ in their characteristics—we design a dual-transformer-based encoder, termed the Multi-Modal Fusion Encoder (MFE). The MFE utilizes a ViT architecture with a class token to semantically align information, serving as the backbone for each modality. Formally, we denote fMRI signals as $B_{f}$, EEG signals as $B_{e}$, and the MFE as $E=\{E_{f},E_{e}\}$. Through the MFE, we obtain:
$f_{f},f_{e},\mathbf{c}_{f},\mathbf{c}_{e}=E(B_{f},B_{e}),$
@@
where $f_{f}$ and $f_{e}$ represent features extracted from fMRI and EEG, respectively, and $\mathbf{c}_{f}$ and $\mathbf{c}_{e}$ are the corresponding class tokens.
Next, we apply a fusion MLP $\psi$ to integrate features from $E_{f}$ and $E_{e}$:
$f_{b}=\psi(f_{f},f_{e}),$
where $f_{b}$ drives the decoder to reconstruct the stimulus.
@@
Simultaneously, we utilize pre-trained contrastive encoders $E_{v}$ and $E_{t}$ to extract embeddings from videos and text generated by the VLM, as illustrated in the modality alignment in Fig. 4. Given video clips with $n$ frames $|V|=\{I_{1},\dots,I_{n}\}$, we first extract embeddings for all frames using $E_{v}$ and then apply an aggregation module $\varphi$ to produce a unified video embedding $\mathbf{c}v$. For text, we directly use the encoder $E_{t}$:
$\mathbf{c}_{v}=\varphi\bigl{(}\{E_{v}(I_{i})\}_{i=1}^{n}\bigr{)}\quad\mathbf{c}_{t}=E_{t}(\text{Text})$
We then compute the contrastive losses as follows:
$\mathcal{L}_{fv}$ $=\mathcal{L}_{clip}(\mathbf{c}_{f},\mathbf{c}_{v});\quad\mathcal{L}_{ft}=\mathcal{L}_{clip}(\mathbf{c}_{f},\mathbf{c}_{t})$
$\mathcal{L}_{ev}$ $=\mathcal{L}_{clip}(\mathbf{c}_{e},\mathbf{c}_{v});\quad\mathcal{L}_{et}=\mathcal{L}_{clip}(\mathbf{c}_{e},\mathbf{c}_{t})$
Therefore, the total contrastive loss for the encoder is defined as:
$\mathcal{L}_{c}=\mathcal{L}_{fv}+\mathcal{L}_{ft}+\mathcal{L}_{ev}+\mathcal{L}_{et}$
During training, we optimize the MFE, including the aggregation module $\varphi$, while keeping the pre-trained image encoder $E_{v}$ and text encoder $E_{t}$ frozen.
After extracting the fused brain feature $f_{b}$, we feed it into the next module for video reconstruction.
@@ 4.2 Neuro Latent Decoder
To effectively reconstruct video stimuli and advance video decoding, our framework leverages CogVideoX [45], which generates videos at 8 fps and 480$\times$720 resolution. Compared to existing approaches [27, 15], our model generates higher-resolution videos with more frames, substantially enhancing brain-based representations. CogVideo employs a causal 3D VAE to encode latent video representations while maintaining temporal causality, thereby preserving the video’s spatiotemporal dynamics.
Specifically, we adapt CogVideo-5B-T2V as the primary diffusion model within our Neuro Latent Decoder (NLD). Given the central role of textual conditioning in video generation models, we replace the original text embedding with our fused brain feature $f_{b}$.
During training, we encode a video clip $V$ using a 3D-VAE $\mathcal{E}$, then introduce noise into the latent representation:
@@
$\mathbf{x}_{0}=\mathcal{E}(V);\quad\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\mathbf{\epsilon}\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
where $t$ is uniformly sampled between 1 and 1000. Following CogVideoX, we adopt explicit uniform sampling of timestamps during training.
As depicted in Fig. 4, we concatenate $\mathbf{x}_{t}$ with the fused brain feature $f_{b}$ for input to the diffusion model. We apply LoRA to fine-tune attention and feed-forward layers within the DiT Blocks of NLD, enabling effective integration of brain-derived signals into video reconstruction
The training objective for NLD is formulated as:
@@
$\mathcal{L}=\mathbb{E}_{\mathbf{V},\mathbf{\epsilon},t}\left|\epsilon-\epsilon_{\theta}\left(\mathbf{x}_{t},\ f_{b},\ t\right)\right|^{2},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
4.3 Extension to Audio Reconstruction
Our framework can readily extend to audio reconstruction. Specifically, to accurately reconstruct audio from brain signals, we leverage the DiT-based audio generation model F5-TTS [5] as our decoder. Since participants simultaneously listened to audio while viewing video stimuli, we fully utilize available audio information. Unlike video reconstruction, we convert pre-segmented audio clips $A$ into text using Whisper-large-v3 [37] for contrastive learning. We extract text embeddings using $E_{t}$, computing contrastive loss exclusively on text modality to avoid visual-audio misalignment noise. Similar to video reconstruction, we apply LoRA to fine-tune attention layers within F5-TTS. Additionally, we adapt the original F5-TTS training strategy by introducing noise to the entire audio segment, aligning with standard diffusion model practices. Further architectural details are provided in the supplementary materials.
5 Experiments
5.1 Cine-Benchmark
Video Metrics. Reconstructing human-centric videos from multimodal brain signals is a challenging task, even within text-to-video frameworks. Thus, establishing effective evaluation metrics is crucial for comprehensively assessing model performance. We evaluate the reconstructed videos on two levels: semantic and perceptual.
1) Semantic-Level: Semantic similarity is essential for evaluating reconstructed videos. Following existing neuro-decoding literature [27, 7, 39, 13, 24], we calculate N-way top-K accuracy across entire videos and individual frames to measure feature similarity. Additionally, we compute the Fréchet Video Distance (FVD) to quantify the distributional similarity between reconstructed and ground-truth videos.
2) Perceptual-Level: Beyond semantics, visual quality is also critical. Therefore, we employ DINO temporal consistency (DTC) [32] and CLIP temporal consistency (CTC) [36] metrics to evaluate video temporal coherence. Additionally, structural similarity (SSIM) and peak signal-to-noise ratio (PSNR) are used to assess frame-level image quality.
Audio Metrics. For audio reconstruction, we similarly adopt metrics at both semantic and perceptual levels:
1) Semantic-Level: Rather than employing the conventional N-way top-K, we utilize Whisper [37] to transcribe the audio into text, enabling the calculation of Word Error Rate (WER) and Character Error Rate (CER) against ground-truth text. We also calculate the Fréchet Audio Distance (FAD) to measure distributional similarity between reconstructed and ground-truth audio signals.
2) Perceptual-Level: We employ Log-Spectral Distance (LSD) and Mel Cepstral Distortion (MCD) to evaluate audio quality by analyzing their spectral characteristics. Additionally, SSIM and Root Mean Squared Error (RMSE) are computed on mel spectrograms to assess perceptual.
5.2 Implementation Details
@@
Video Reconstruction. We employ the AdamW optimizer with $\beta=(0.9,0.95)$ and an initial learning rate of $1\times 10^{-4}$. Due to GPU memory constraints, CineSync is trained using a two-stage approach. In the first stage, only the multimodal fusion encoder is trained using contrastive loss with a batch size of 16 on a single H100 GPU. To enhance the effectiveness of contrastive learning, we augment the training dataset with diverse textual captions generated by Qwen2.5-VL-7B [2] and Llava-Video-7B [46]. This pretraining phase lasts approximately 50 epochs. In the second stage, the pretrained encoder is integrated into the full model, which is finetuned for 5000 steps on four H100 GPUs using a batch size of 2 per GPU. The fMRI and EEG transformers each consist of 24 layers with a transformer dimension of 2048 and a token length of 227 (226 spatial tokens plus one class token). The LoRA configuration uses a rank of 128 and a scaling factor $\alpha$=64. Input data consist of 4-second multimodal brain signals, yielding standard inputs of 5$\times$8405 fMRI voxels and 64$\times$4000 EEG data points.
Audio Reconstruction. We use the AdamW optimizer with $\beta$=(0.9, 0.95) and an initial learning rate of 1$\times$10-4. Given the strict alignment between text and audio , we reduce the vision feature and utilize transcriptions generated by Whisper [37] to facilitate contrastive learning. We use F5-TTS-Base [5] as the diffusion model in the neural latent decoder and train the model for 30,000 steps using a batch size of 8 on a single A6000 GPU. Both fMRI and EEG transformers comprise 12 layers, each with a transformer dimension of 1920 and a token length of 377 (376 spatial tokens plus one class token). The LoRA configuration uses a rank of 64 and a scaling factor $\alpha$=32. Similar to video reconstruction, each input contains standard 4-second segments, resulting in 5 $\times$ 10541 fMRI voxels and 64 $\times$ 4000 EEG data points.
5.3 Results of Video Reconstruction
As the first attempt to combine fMRI and EEG signals for video decoding, there are no directly comparable baselines. To validate the effectiveness of our proposed CineSync and assess the quality of the CineBrain dataset, we select two baselines: CineSync-fMRI and CineSync-EEG, which remove the EEG or fMRI branches from our model, respectively. Additionally, we include GLFA [27], a strong end-to-end baseline based solely on fMRI signals. All models are trained and evaluated on the same dataset. Tab. 2 presents the averaged metrics across all subjects. The results demonstrate that CineSync-fMRI significantly outperforms GLFA across all metrics by a substantial margin, confirming the effectiveness of our framework. Moreover, CineSync surpasses both CineSync-fMRI and CineSync-EEG, clearly indicating that our framework successfully integrates fMRI and EEG signals to enhance video reconstruction quality. This demonstrates the superiority of CineSync and establishes a new state-of-the-art in reconstructing videos from human brain signals. Additionally, qualitative results of subject 3, shown in Fig. 5, indicate that CineSync captures more temporal details compared to the two baselines, demonstrating that our framework effectively leverages EEG’s temporal resolution to complement fMRI’s spatial characteristics. Furthermore, the experimental results affirm the high quality of our dataset, demonstrating its value in supporting multimodal brain-signal research. We hope our dataset will facilitate further advancements in computer vision and cognitive neuroscience.
5.4 Results of Audio Reconstruction
To verify the flexibility and robustness of our framework, we also conduct experiments reconstructing audio signals from multimodal brain data. In this setting, we again select CineSync-fMRI and CineSync-EEG as baselines due to their strong performance in video reconstruction. Qualitative and quantitative results are provided in the supplementary material. These results indicate that our framework can be easily adapted to audio reconstruction tasks, further validating its effectiveness and the high quality of our dataset.
5.5 Exploration of Brain ROIs
As discussed in the introduction, different brain regions work collaboratively. Based on this insight, we conducted an exploratory study to investigate the effectiveness of joint brain-region interactions. Specifically, we incorporated auditory ROIs into the fMRI signals for video reconstruction tasks and visual ROIs into the fMRI signals for audio reconstruction tasks while retaining all the EEG data, resulting in a total of 18,946 voxels per frame. We refer to this modified model as CineSync. As shown in Tab. 2 and the supplementary material, CineSync achieves better results than CineSync in both video and audio reconstruction. These experimental findings demonstrate that utilizing jointly stimulated brain regions can effectively enhance the performance of video and audio reconstruction from brain signals.
@@ 6 Conclusion
In this paper, we present the first large-scale multimodal audiovisual stimulation dataset, CineBrain, supporting various reconstruction tasks. Building on this dataset, we develop an innovative framework, CineSync, successfully fuses fMRI and EEG signals by leveraging the high spatial resolution of fMRI and the high temporal resolution of EEG to enhance both video and audio decoding performance. In the first stage, the multimodal fusion encoder extracts features from both fMRI and EEG signals and aligns them with visual and textual representations at the semantic level. In the second stage, the framework employs a neuro latent decoder, wherein a LoRA-tuned diffusion decoder reconstructs the stimuli based on these extracted multimodal features. We also introduce Cine-benchmark to evaluate our model’s performance in both video and audio reconstruction. CineSync achieves SOTA performance in video reconstruction and demonstrates sufficient flexibility and robustness for audio reconstruction. Our experiments empirically validate the effectiveness of our dataset, highlighting its potential to support future research. We hope that our framework will serve as a strong baseline for multimodal brain decoding and that our dataset will pave the way for further advancements in computer vision and cognitive science.
Appendix A Additional Experimental Results
A.1 More Results of Video Reconstruction
To demonstrate the quality and temporal consistency of the reconstructed videos generated by our method, we present additional examples containing 12 frames each in Figures B.2, B.3, and B.4. Overall, our reconstructions exhibit strong semantic and perceptual performance.
A.2 Architecture of CineSync for Audio
The detailed architecture of CineSync, employed for the audio reconstruction task, is depicted in Fig. A.1. To simplify the illustration, we omit the contrastive loss of the Multi-Modal Fusion Encoder and focus on the differences from the video reconstruction pipeline. Specifically, we first convert an audio file into a mel spectrogram. We then use Vocos [40] as encoder $\mathcal{V_{E}}$ to obtain a latent representation of the mel spectrogram $M$, and subsequently add noise to it for training:
$\mathbf{x}_{0}=\mathcal{V_{E}}(M);\quad\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\,\mathbf{\epsilon},\quad\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
Then, we concatenate the fused brain features with $\mathbf{x}_{t}$ to form the training input for our neural latent decoder. Unlike the strategy used for the video decoder, we train the neural latent decoder using a flow-matching approach.
A.3 Results of Audio Reconstruction
The quantitative results, averaged across all participants, are summarized in Table B.1. Our CineSync approach outperforms both baselines, indicating that our framework effectively leverages the strengths of fMRI and EEG signals for audio reconstruction. Additionally, qualitative comparisons between CineSync and the baseline methods are shown in Figure B.5. From the mel spectrograms, it is evident that CineSync achieves superior performance compared to baselines relying on a single modality. These experimental results in audio reconstruction further demonstrate that integrating fMRI and EEG improves performance, consistent with our findings on video reconstruction. Furthermore, we also explored incorporating fMRI voxels from visual areas into the audio reconstruction task, which similarly improved reconstruction quality. This outcome suggests that different brain regions’ activations are interrelated. Overall, our experiments empirically validate the high quality of the CineBrain dataset, which can support various experimental settings.
Figure A.1: Architecture of CineSync for audio reconstruction. We concatenate the fused brain features with the noised latent representation of the mel spectrogram as the training input. We then apply LoRA [21] to the attention and feed-forward layers in the DiT [34] blocks of our neural latent decoder.
Appendix B EEG Experimental Device

### v2

@@ 4 Methods
Table 2: Performance comparison of different multimodal encoder structures. We evaluate all variants using video-level semantic metrics and frame-level perceptual metrics. Bold denotes the best performance, while underlined denotes the second-best.
Methods Semantic-level Perceptual-level
[…]
(a) Joint Transformer 0.924 0.274 128.0 0.232 9.30
(b) Two-Stage Fusion 0.921 0.278 120.0 0.242 9.60
(c) Cross-Attn. Fusion 0.921 0.292 110.0 0.249 9.90
(d) SpatialCat (f113-e113) 0.918 0.276 114.0 0.255 9.95
(d) SpatialCat (f34-e192) 0.928 0.311 106.0 0.250 10.10
(d) SpatialCat (f24-e202) 0.926 0.307 108.0 0.247 10.00
(e) Dual Trans. Fusion 0.929 0.324 51.53 0.249 12.03
@@ 4.1 Multi-Modal Fusion Encoder
fMRI–EEG Fusion Exploration. As an initial exploration of fMRI–EEG integration, we evaluate five fusion architectures under comparable parameter budgets, as illustrated in Fig. 4. (a) Joint Transformer: concatenates fMRI and EEG features and feeds them into a unified Transformer for joint representation learning. (b) Two-Stage Fusion: processes fMRI and EEG separately with individual Transformers, followed by a joint Transformer for integration. (c) Cross-Attention Fusion: employs modality-specific Transformers where each block includes cross-attention for inter-modal information exchange. (d) Spatial Concatenation: uses different token counts for fMRI and EEG features, doubles the feature dimension, and spatially concatenates them before fusion. For a fair comparison, the total token count is halved. (e) Dual Transformer Fusion: extracts features from each modality with independent Transformers and aggregates them at the final stage. Tab. 2 indicates that separately encoding fMRI and EEG signals yields better performance. This finding reveals a substantial discrepancy between the two modalities, suggesting that directly applying fully shared self-attention is not suitable for their fusion. It also provides important insights that inform the design of our subsequent fusion module.
@@
Motivated by this observation, we design a dual-transformer architecture termed the Multi-Modal Fusion Encoder (MFE). Inspired by the success of contrastive learning [clip] in neuro-decoding [mind-video, lichong_video, mind3d++] and multi-modal representation learning, the MFE adopts the Vision Transformer (ViT) as the backbone with a class token to semantically align information across modalities. Formally, we denote the fMRI and EEG signals as $\mathbf{x}_{f}$ and $\mathbf{x}_{e}$, respectively. The MFE is defined as $E=\{E_{f},E_{e}\}$, where $E_{f}$ and $E_{e}$ are modality-specific encoders for fMRI and EEG. Through the MFE, we obtain both latent features and class tokens:
$\mathbf{z}_{f},\mathbf{z}_{e},\mathbf{c}_{f},\mathbf{c}_{e}=E(\mathbf{x}_{f},\mathbf{x}_{e}),$ (1)
@@
where $\mathbf{z}_{f}$ and $\mathbf{z}_{e}$ denote the latent representations extracted from fMRI and EEG, and $\mathbf{c}_{f}$ and $\mathbf{c}_{e}$ are their corresponding class tokens used for contrastive alignment.
To integrate information across modalities, we introduce a fusion MLP $\psi$ that combines the latent features:
$\mathbf{z}_{b}=\psi(\mathbf{z}_{f},\mathbf{z}_{e}),$ (2)
where $\mathbf{z}_{b}$ serves as the unified brain representation for subsequent video reconstruction.
@@
For cross-modal alignment, we employ pre-trained contrastive encoders $E_{v}$ and $E_{t}$ to extract embeddings from videos and their textual descriptions generated by the VLM (see Fig. 5). Given a video clip with $n$ frames $\mathcal{V}=\{I_{1},\dots,I_{n}\}$, frame-wise embeddings are obtained via $E_{v}$ and aggregated by a temporal aggregation module $\varphi$ to form a video-level representation $\mathbf{c}_{v}$. For text, we directly encode the caption using $E_{t}$:
$\mathbf{c}_{v}=\varphi\bigl(\{E_{v}(I_{i})\}_{i=1}^{n}\bigr),\quad\mathbf{c}_{t}=E_{t}(\text{Text}).$ (3)
We define the contrastive objectives as:
$\mathcal{L}_{fv}$ $=\mathcal{L}_{\text{clip}}(\mathbf{c}_{f},\mathbf{c}_{v}),\quad\mathcal{L}_{ft}=\mathcal{L}_{\text{clip}}(\mathbf{c}_{f},\mathbf{c}_{t}),$ (4)
$\mathcal{L}_{ev}$ $=\mathcal{L}_{\text{clip}}(\mathbf{c}_{e},\mathbf{c}_{v}),\quad\mathcal{L}_{et}=\mathcal{L}_{\text{clip}}(\mathbf{c}_{e},\mathbf{c}_{t}).$
To further align fMRI and EEG representations, we introduce an additional cross-modal contrastive loss:
$\mathcal{L}_{fe}=\mathcal{L}_{\text{clip}}(\mathbf{c}_{f},\mathbf{c}_{e}).$ (5)
Thus, the overall contrastive objective is given by:
$\mathcal{L}_{\text{c}}=\mathcal{L}_{fv}+\mathcal{L}_{ft}+\mathcal{L}_{ev}+\mathcal{L}_{et}+\mathcal{L}_{fe}.$ (6)
During training, we optimize the MFE, the fusion MLP $\psi$, and the aggregation module $\varphi$, while keeping the pre-trained encoders $E_{v}$ and $E_{t}$ frozen. After extracting the fused brain representation $\mathbf{z}_{b}$, it is fed into the subsequent module for video reconstruction.
@@ 4.2 Neuro Latent Decoder
To effectively reconstruct the temporal stimuli, we design a Neuro Latent Decoder (NLD), a general-purpose module that can be seamlessly integrated into any video diffusion model. Specifically, for video reconstruction, we adopt and adapt CogVideoX-5B [cogvideox] as the underlying diffusion model within our NLD. CogVideoX-5B generates videos at 8 fps with a resolution of $480\times 720$, conditioned on text prompts. Given the central role of textual embeddings in video diffusion models, we replace the original text condition with our fused brain representation $\mathbf{z}_{b}$.
During training, a video clip $V$ is first encoded into the latent space using a 3D-VAE $\mathcal{E}$, after which noise is added according to the forward diffusion process:
@@
$\mathbf{x}_{0}=\mathcal{E}(V),\quad\mathbf{x}_{t}=\sqrt{\overline{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\overline{\alpha}_{t}}\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (7)
where $t$ is uniformly sampled from $\{1,\dots,1000\}$. Following the training strategy of CogVideoX, we adopt explicit uniform timestep sampling. As illustrated in Fig. 5, the noised latent $\mathbf{x}_{t}$ is concatenated with the fused brain feature $\mathbf{z}_{b}$ and then fed into the diffusion model.
To efficiently adapt the model to brain-derived inputs, we employ LoRA fine-tuning on both the attention and feed-forward layers within the DiT blocks of NLD, enabling effective integration of neural representations into the video generation process. The training objective of the NLD follows the standard diffusion loss:
@@
$\mathcal{L}=\mathbb{E}_{V,\,\boldsymbol{\epsilon},\,t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},\,\mathbf{z}_{b},\,t)\right\|^{2}\right],\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (8)
@@
GLFA [lichong_video] 0.801 0.167 128.76 0.847 0.225 0.706 0.735 0.123 7.526
NeuroClips [gong2024neuroclips] 0.816 0.183 116.36 0.833 0.142 0.871 0.868 0.087 6.854
@@
5 Experiments
Metrics. To comprehensively evaluate the model’s performance, we assess the reconstructed videos at two levels: semantic and perceptual. 1) Semantic-Level: Semantic similarity is essential for evaluating reconstruction quality. Following prior neuro-decoding studies [lichong_video, mind-video, mind-eye, mind3d, eeg2speech], we compute N-way top-K accuracy across entire videos and individual frames to measure feature similarity. Additionally, the Fréchet Video Distance (FVD) is calculated to quantify the distributional similarity between reconstructed and ground-truth videos. 2) Perceptual-Level: Beyond semantics, visual quality is also critical. We employ DINO temporal consistency (DTC) [oquab2023dinov2] and CLIP temporal consistency (CTC) [clip] to assess video temporal coherence. Structural similarity (SSIM) and peak signal-to-noise ratio (PSNR) are further used to evaluate frame-level visual fidelity.
5.1 Implementation Details
Both the fMRI and EEG transformers contain 12 layers with a hidden dimension of 2048 and a token length of 227 (226 spatial tokens plus one class token). The LoRA configuration in the NLD adopts a rank of 64 and a scaling factor $\alpha=64$. For modality alignment, we employ SigLIP [siglip] to extract visual and textual embeddings. Each input sample comprises 4-second multimodal brain signals, corresponding to $5\times 8{,}405$ fMRI voxels and $64\times 4{,}000$ EEG data points. We use the AdamW optimizer with $\beta=(0.9,0.95)$ and an initial learning rate of $1\times 10^{-4}$. Further implementation details are provided in the supplementary material.
5.2 Experimental Results
Encoder Comparison. As shown in Tab. 2, we observe that reducing the interaction between fMRI and EEG during the feature extraction stage leads to better overall performance. This finding suggests that the two modalities exhibit substantial representational differences, and excessive early fusion may hinder effective feature learning. Interestingly, experiment (d) shows that when the total token count is fixed at 226, moderately increasing the number of EEG tokens leads to performance gains in video reconstruction. However, we also observe that although this design expands the feature dimension, the reduced overall token count yields suboptimal performance compared with method (e), indicating that the number of tokens plays a more critical role.
Comparison with Baselines. To evaluate the effectiveness of our proposed CineSync, we compare it with several representative baselines, including EEG2Video [eeg2video], GLFA [lichong_video], and NeuroClips [gong2024neuroclips], as well as two simplified variants of our model: CineSync-fMRI and CineSync-EEG. All models are trained and evaluated on the same dataset for fair comparison. Tab. 3 reports the average performance across all subjects. Overall, the two simplified versions of CineSync substantially outperform the baselines across all metrics, validating the robustness of our framework even when using a single modality. Furthermore, the full CineSync model surpasses both CineSync-fMRI and CineSync-EEG, demonstrating that jointly modeling fMRI and EEG signals enables more comprehensive decoding of brain activity and leads to higher-quality video reconstruction. These results establish CineSync as a new state-of-the-art framework for reconstructing dynamic visual stimuli from human brain signals. In addition, we conduct an extended experiment by incorporating auditory ROIs into the fMRI inputs for video reconstruction, while retaining all EEG data, resulting in a total of 18,946 voxels per frame. We refer to this variant as CineSync. The results show that incorporating auditory ROIs improves video reconstruction performance, underscoring the benefits of multimodal and cross-regional integration in brain decoding.
Qualitative Analysis. Qualitative results in Fig. 6 confirm that CineSync captures finer temporal dynamics and richer perceptual details than the baselines. This suggests that our framework effectively leverages the complementary advantages of fMRI and EEG to enhance reconstruction fidelity. Moreover, the consistently high reconstruction quality across subjects highlights the reliability and value of the CineBrain dataset in supporting multimodal brain-signal research, potentially facilitating further progress at the intersection of computer vision and cognitive neuroscience.
Ablation Studies. We conduct an ablation study to evaluate the effectiveness of the contrastive learning–based alignment employed in our framework. The results, presented in the supplementary material, demonstrate that our multi-level contrastive alignment strategy significantly contributes to the overall performance, confirming its effectiveness in aligning multimodal brain representations.
@@ 6 Conclusion
In this paper, we present CineBrain, the first large-scale multimodal dataset simultaneously recording fMRI and EEG under naturalistic audiovisual stimulation. Building on this dataset, we introduce the task of reconstructing continuous video stimuli from multimodal brain signals. Furthermore, we explore effective strategies for integrating fMRI and EEG and propose CineSync, an innovative framework that integrates fMRI and EEG to enhance video reconstruction quality. CineSync achieves state-of-the-art performance, consistently outperforming baselines using single-modal input, thereby demonstrating the benefits of multimodal brain representations. Experiments further show that incorporating auditory ROIs improves visual decoding, reinforcing the reliability and value of CineBrain dataset and highlighting its potential to advance multimodal brain decoding research. We hope our dataset and framework will serve as strong baselines and inspire future progress at the intersection of computer vision and cognitive neuroscience.
[…]
Supplementary Material
7 Detailed Information on CineBrain
7.1 Detailed ROIs Selected in the Experiments
The regions of interest (ROIs) used in our study are summarized below. For the visual cortex, we include the following ROIs: V1, V2, V3, V3A, V3B, V3CD, V4, LO1, LO2, LO3, PIT, V4t, V6, V6A, V7, V8, PH, FFC, IP0, MT, MST, FST, VVC, VMV1, VMV2, VMV3, PHA1, PHA2, PHA3. For the auditory-related cortex, we select ROIs corresponding to: 4, 7AL, 7Am, 7m, 7PC, 7PL, 7Pm, 8Ad, 8Av, 8BM, 8C, 9a, 9p, 10d, 10v, 44, 45, 46, 47m, A1, IPS1, p32, PGp, s32, STGa, STSda, STSdp, STSva, STSvp, TPOJ1, TPOJ2, TPOJ3. In total, the selected visual ROIs include 8,405 voxels, while the auditory ROIs include 10,541 voxels.
7.2 Prompts for Generating Text Descriptions
Textual descriptions play an important role in our training pipeline. The prompts used to generate these descriptions with Qwen2.5-VL [Qwen2.5-VL] are provided in Fig. 7.
7.3 EEG Experimental Device